OpenStudy (biue):

What is 7(n+4) = 21 equal to?

geerky42 (geerky42):

Solve for n? Divide both sides by 7, then subtract both sides by 4.

OpenStudy (biue):

7(n+4) = 21 Choices. n = -1 n = 7 n = 10 n = 1 Im having trouble :/

OpenStudy (biue):

@geerky42

geerky42 (geerky42):

You want to isolate n. So first you need to "remove" 7. w can do that by divide it by 7, but we must do that to 21 too. \[7(n+4)=21\quad\Longrightarrow\quad\dfrac{7(n+4)}7=\dfrac{21}7\]

geerky42 (geerky42):

Then we are left with \(n+4 = 3\).

OpenStudy (biue):

Ok.

OpenStudy (biue):

So n = 7? Im not really good with mathematics sorry.

geerky42 (geerky42):

n is still not isolated, so we need to remove 4. Opposite of adding 4 is subtracting 4. So we subtract \(n+4\) by 4, do the same to 3. \[n+4=3\quad\Longrightarrow\quad n+4-4=3-4\] \[\boxed{n=-1}\]

OpenStudy (biue):

Thank you.